Volume Imaging Lidar Operations Record at Lake-ICE

This page contains a brief record of VIL operations during Lake-ICE.
Return to the Index. Return to Current Operations Menu
Check out Status of VIL at Lake-ICE here.

Lake-Induced Convection Experiment, Sheboygan, WI


Summary: Day two of cold air outbreak. Skies cleared yesterday soon after we shutdown (about noon). Evening temperatures dropped into the teens but were about 20F by 6 am CST today. Also, skies were overcast by dawn this morning. Early visibilities were reduced due to haze (not steam or fog) and improved gradually by mid-morning.

This mornings atmosphere was often complicated by clouds and streaks of falling snow. A shallow mixing layer (100-200 m thick) was observed most of the time and appeared to be decoupled from the overlying cloud deck.

Later in the morning the precipitation stopped and a ML with sharp entrainment zone formed. At 18 Z the ML top was at about 400-500 m AGL and there was a cloud deck at 3 km with streaks of precipitation falling down to 2 km.

RHI Volume scans: 18:54-19:06 Z; 90-130 deg azi, 0-15 ele.

RHI Volume scans: 17:47-18:52 Z; 75-105 deg azi, 0-15 ele. Start of side disk 7 side A.

At 17 Z a hole in the cloud deck appeared over us. Very large. Another one was spotted to our NNW. Looks like cirrus from above seeded a supercooled cloud deck and created a hole. Ralph took pictures.

RHI Volume scans: 16:32 til about 17:40 Z; 75-105 deg azi, 0-15 ele. Stopped to change disk. End of 6A. Start of 7A. Forgot to start 6b.

SGNW3 observation from 16 Z: -4.1C (24.6F), dp=-6.1C (21F), wind: from 300 deg at 2 m/s. PPI scans: 15:30-16:00 Z; 70-176 deg azi, ele=1.5 deg.

RHI scans: 14:57-15:27 Z; 120 deg azi; 0-15 ele.

RHI scans: 14:31-14:56 Z; 176 degrees azimuth; 0-15 degrees elevation.

PPI scans: 13:52 - 14:31 Z ; 70-176 degrees azimuth, .05 degrees elevation.

SGNW3 observation at 13Z: -4.7C, dp=-6.5C, wind: 280 @ 5m/s Note: `hazy conditions' this morning. No fog or steam. SDM

RHI volume scans: 12:24 - 13:49 Z (azimuth: 146-176, elevation: 0-15.)


Summary: Day one of another cold air outbreak. This time not as cold as previous cases. Lots of clouds and flurries early. Mixed layer very shallow.

RHI volume scans: 17:04 - 18:05 Z (azimuth: 90-130, elevation: 0-15.)

RHI scans: 16:35 - 16:53 Z (azimuth: 110, elevation 0-15.)

PPI scans: 15:46 - 16:21 Z (azimuth: 70-176, elevation: .05)

RHI Volume scan: 12:47-15:21 (6:47-7:21 am CST) 156-176 degrees azimuth, 0-15 degrees elevation.

PPI scan: 12:33-12:43 Z (6:33-6:43 am CST), 70-176 degree azimuth, 1 degree elevation.

RHI Volume scan: 11:20-12:22 Z (5:20-6:22 am CST), 80-120 degrees azimuth, 0-15 degrees elevation.

RHI Volume scan: 10:49-11:12 Z (4:49-5:12 am CST), 70-110 degrees azimuth, 0-15 degrees elevation.


Summary: Day one of another excellent cold-air-outbreak. Dawn temperatures near zero with strong westerly flow at surface. Clear skies over land and at coast with dense steam-fog over the lake and a cloud or fog bank constantly offshore.

The Wyoming King air did a series of north-south stacked legs at the longitude of the buoy this morning from 19:12 - 19:55 Z. The legs were about 20 km long and centered on the buoy. The altitudes of the legs were 1000, 1300, 1600, and 2200 ft AMSL (The surface of Lake Michigan is about 700 ft AMSL). The last VIL data collected before the stacked-legs began was at 18:37 Z.

12:47-13:37 Z 40-degree wide RHI volume scans pointing downwind (110 degrees). [12 Z observation: -4.7 F, wind 5 m/s from 280. Dew point -13.4 F.] 400 m thick BL and TIBL observed. At 13:37 the sun was severely affecting our data and we decided to switch to PPI scans to avoid this problem. However at that time we discovered problems with the laser water temperature and were not able to resume data collection until 15:48 Z.

15:48-16:22 Z 20-degree wide RHI volume scans pointing downwind.

16:28-16:32 Z wide angle PPIs (70-176 degrees at 0 degree elevation). These data did not look very useful so we stopped and tried to start downwind RHIs, but the BSU failed.

16:44-17:08 Z Downwind RHIs

17:19-18:37 downwind RHI volume scans. Azimuth range from 90-130 degrees. Elevation 0-15 degrees.

NOTE #1: Chiller failed this morning because it got too cold. We temporarily fixed it by bypassing the cold-protection relays (i.e. jamming a screwdriver between the relay and the tank to hold the relay closed.) Jim says if we had left the shrouding all closed on the chiller it wouldn't of happened.

NOTE #2: Beam steering unit failed several times this morning. Note sure of cause. Ralph discovered power cord to the data system loose in the socket on the wall. Warm air was blown up into the BSU and an IR heater was placed on the roof to warm it from the outside. The BSU failures occurred only when we went to `go home' between scan patterns. At that point we were forced to reset everything including power to the BSU and restarting VIL control software. Once we were forced to manually move the BSU back to the home position before we could control it by software.


Summary: Day two of an excellent cold-air-outbreak. Dawn temperatures near 0F with light westerly flow at surface. Mostly clear skies over land and along the shore. Thick steam fog in shallow layer over lake with clouds forming about 10 km offshore.

1/11/98 9:45 am CST. No data collected from 14:30-15:30 Z (8:30-9:30 CST) due to laser cooling problem. Resumed data collection 15:30 Z (9:30 CST). TIBL no longer obvious in RHI scans. ABL height about 200 m AGL beyond 1.5 km offshore--slightly higher near shore. We are collecting RHI volume scans from 70-110 degrees azimuth and 0-15 elevation. (14 Z air temperature 0.3 F, wind W at 6 m/s. skies mostly clear overhead, steam fog over lake, offshore clouds starting at about 10 km range.)

1/11/98 7:49 am CST. Resumed data collection after repairing laser cooling system. Collecting RHI volume scans from 70 - 110 degrees azimuth (20 degrees on both side of due east) and 0 - 15 degrees elevation. TIBL still present. Steam fog beautiful. 13Z air temperature 0.7 F (-17.4 C), wind from 260 degrees at 5 m/s.

1/11/98 7:00 am CST. Arrived extra early for ideal conditions (5:20 AM CST). Observed thermal internal boundary layer (TIBL) over lake from 5:51-6:16 am until cooling system on laser failed. Working on repair and expect to be operational within hour.


Day one of an excellent cold-air-outbreak. Dawn temperatures slightly above 0 with strong southwesterly flow.


  • Partly cloudy, Thin, broken cu moving to the Northwest. Scattered ci. Sunrise temperature approximately 25 deg F. Surface winds light and variable with slight offshore component. Lake surface is calm with wind line visible offshore. Power plant plumes pointing to the Northwest in moderate breeze. Lidar images show strong land breeze front 1-3 km offshore.

  • 12:43-13:11 UTC RHI volume scan. Az=85-135 deg, Elev=0-15 deg, d_az=0.5, d_elev=0.23

  • 13:12-15:21 UTC, RHI volume scan. Az=126-176 deg, Elev=0-15 deg, d_az=0.5, d_elev=0.23 deg.

  • 15:24-16:46 UTC, PPI scan, Az=85-176 deg, Elev=0 deg, d_az=0.08 deg. Extermely light snow shower approaches from Southeast and passes through scan area near end of period. Snow is easily seen to end of data range at 18 km. Land breeze front is clearly recorded to about 15 km South of lidar.

  • 16:51-17:12 UTC. RHI scan pointing directly into cloud layer wind. Elev=0-15 deg, Az=165 deg, d_elev=0.23 deg. Land breeze front begins to breakdown at begining of this observation period and aerosol marking front is rapidly blown ashore. Water surface becomes rough and waves build to approximately 1 foot.


  • Very light snow showers over lake--snow at lidar van was barely visible as scattered snow grains which we did not notice until we saw the lidar display. Visibility was excellent with lidar structure visible to 18 km on nearly all scans.

  • 13:25-15:03 UTC RHI volume scan, az=85-135 deg, elev=0.25-15 deg, Cloud base at 800 m. Convective plumes with tops at 100 m.

  • 15:18-16:39 UTC PPI scan. Az=70-175 deg, elev=0.25, d_Az=0.08 deg.

  • 16:43-17:28 UTC RHI volume scan, Az=70-175 deg,elev=0.25-15 deg d_elev=0.23, d_az=0.5 deg.

  • 17:43-17:57 UTC RHI scans into lake surface at several azimuths to check system leveling.

  • 19:18-20:36 UTC RHI volume scan. Az=85-135 deg, elev=0.25-15,d_az=0.5, d_elev=0.23 deg.

  • 20:55-21:57 UTC PPI scan. Az=85-176 deg, Elev=0.0 deg, d_az=0.08 deg.

  • 22:30-22:46 UTC, RHI volume scan, Az=70-175 deg, Elev=0.25-15 deg, d_az=0.5, d_elev=0.23 deg.

  • 22:57-23:16 UTC, PPI scan at increased elevation to allow scanning over shore line to the south. Az=120-210 deg, Elev=2.0 deg, d_az=0.08 deg.


  • 14:00 UTC Water temperature measured at shore was 35 deg F and air temperatue at lidar van was 25 deg F. Light westerly surface wind. Power plant plumes show easterly wind aloft.

  • 14:31-15:07 UTC RHI volume scan, az= 60-120 deg, elev=.25-15 deg., Mostly clear sky, shallow convection.

  • 15:19-15:54 UTC RHI volume scan, az=70-110 deg, elev=.25-15 deg.

  • 15:59-16:32 RHI scans to find lake surface in order check leveling of system.

  • 17:58-18:27 UTC PPI Az=80-175 deg. elev.=0.25 deg, Excellent small scale structure thoughout scan. Sky mostly clear, excellent visibility. Light winds.

  • 18:28-19:00 UTC RHI volume scan, azi=70-110 deg, elev.= 0.25-15 deg., Well defined boundary layer with tops at approximately 100 m.

  • 19:03-19:18 UTC RHI volume scan, azi=85-180 deg, elev.=0.25-10 deg.

  • 19:23-20:02 UTC PPI azi=75-175 deg, elev=0.25 degrees. Surface wind indicated by flags on shore is southeast while power plant plumes indicate south west winds aloft. Excellent structure throughout scan.

  • 20:20 Water temperature at shore 39 deg F air temperature at lidar van 41 deg F.

  • 12/13/97

    No operations.



    No operations.


  • 20:46-21:10 UTC RHI volume scan--system testing.

  • 23:45-23:54 UTC RTI vertical pointing--system testing.

  • 12/09/97

    No operations.


    No operations.


    No operations.


    No operations.


  • 15:32-15:53 UTC, RHI volume scan, elev. 0.5-15, dphi=0.23. Light snow, complex aerosol and cloud structure, Max range limited by attenuation to approx. 7 km.

  • 16:20-17:03 UTC, RHI volume scan, elev. 0.5-15. dphi=0.23. Conditions same as previous.

  • 17:50 UTC PPI scan, elev=2.0, Navigation scan across power plant stacks.

  • 20:01-20:14 UTC RHI volume scan, Az. 65-115 deg, dtheta=1.0, elev=0.5-15 deg, dphi=0.23. Scattered Cu bases 300 to 500 meters. Range limited by attenuation to approx. 8 km. Convective plume tops at 300 m

  • Return to the Index. Return to Lake-ICE

    UW Lidar // Dec 2, 1997 // root@lidar.ssec.wisc.edu